# What is emergence, after all?

### v2

@@ A New Ontology
The emergence of the giant component is foundational: the Internet stays functional despite random failures because its giant component remains intact (70). Likewise, in epidemic dynamics, if an infected person enters a social network and contacts someone in the giant component, the disease can potentially spread to nearly everyone after enough time. In this way, the onset of an epidemic reflects the same percolation principles that govern network connectivity (71, 72). We say that an outbreak occurs or an epidemic emerges when a significant number of individuals in a population become infected. The onset, size, and time scales of this emergence can be measured and modeled through the contagiousness of the disease and how the contact network is structured and changes over time (73). The emergence of large-scale connectivity is one example among many in networks. Furthermore, it is also easy to argue that communities emerge when networks evolve with local rules (74).
Emergence of Herd Immunity
@@
Vaccination is a key strategy for suppressing epidemics. Take measles, a highly contagious disease with a basic reproduction number of $R_{0}\sim 15$—meaning one case can infect around fifteen others early on in a fully susceptible population (75, 76). Before routine vaccination in the UK began in the late 1960s, outbreaks recurred every 2–3 years as the number of susceptible children quietly rose above the transmission threshold.
@@
Vaccination protects in two ways: directly, by immunizing individuals and removing them as nodes from the transmission network, and indirectly, as Fig. 3(a-c) shows, by forming a firewall (or firebreak (77)) and breaking the chains along which the infection might spread (78). In a vaccinated population, before the introduction of a new infection, individuals are either susceptible ($\mathcal{S}$) or immune ($\mathcal{R}$). The network representation of this population can always be rearranged to separate these two groups, with the interface formed by links connecting $\mathcal{S}$ and $\mathcal{R}$ nodes (Fig. 3f). The size of the immune set represents the direct benefit of vaccination, while the number of interface links serves as a proxy for indirect protection.
@@
Recent studies formalize this mechanism using bond percolation and message-passing on real and synthetic networks (81, 79). They track the dynamics of the susceptible giant component and the firefront density, $\rho_{\mathcal{SR}}$, across diverse mixing patterns. These analyses reveal a competing mechanism: targeting superspreaders or implementing acquaintance immunization increases collective immunity, while the spatial structure of the network can hinder it by localizing the protection, leaving remote, susceptible pockets vulnerable (81).
Herd immunity represents a paradigmatic case of emergence in social networks; Local immunizations combine nonlinearly to produce a global shield shaped by the network’s structure. Recognizing this not only deepens our understanding of collective behavior but also equips us with practical tools to design smarter, more efficient epidemic interventions.
@@
In scientific practice, “emergence” almost always refers to weak emergence: a phenomenon grounded in the system’s components and interactions at the right scale, with no need for non-physical explanations. One can argue that a microscopic theory is incomplete or inapplicable to some macroscopic regimes—but to invoke strong emergence in domains like consciousness or social behavior (84, 85) amounts to introducing new causal principles not anchored in substrate dynamics, and thus steps outside the scope of scientific method (86). Therefore, if a macro-level description effectively explains a phenomenon—even more parsimoniously, with fewer micro-level details than a microscopic one—nothing beyond physicalism is at play. There have been, of course, historical movements that made strong-sounding claims about emergent phenomena without yielding testable or predictive frameworks (87, 88, 89, 90, 91, 92).
As causality connects events within the same scale, emergence connects descriptions across scales, revealing how collective patterns arise from local interactions. But the influence does not only run upward. Much like the walls of a container shape the motion of the particles inside, higher-level structures can constrain or guide the behavior of their components without violating the underlying laws. This interplay is often described as downward or top-down causation (93, 94).
It is tempting to wonder whether complex aesthetic outcomes might also emerge from simple, knowable constraints—a question deeply resonant with the concept of emergence itself. If Poe’s self-portrait in The Philosophy of Composition (1846) is taken at face value, at least one kind of aesthetic complexity is not strongly emergent; it can be reconstructed from an explicit, step-wise recipe. Poe famously crafted the poem The Raven by first fixing a desired emotional effect—melancholy—and then deliberately selecting every element (theme, meter, and the refrain “Nevermore”) to achieve that goal (95). He insists good writing is methodical and analytical, explicitly rejecting artistic intuition. While many, including T. S. Eliot, argue that something ineffable lies beyond algorithmic methods (96), Poe’s essay reminds us that not all creativity is opaque to analysis—even if many artists and readers prefer it that way. Today, sentiment-analysis tools can map—and perhaps replicate—a narrative’s emotional arc (97, 98), sharpening our insight into how writers engineer intricate effects. However, until we can define such processes rigorously and run controlled experiments on well-defined creative variables, any final judgment about whether art is ultimately rule-bound or irreducibly spontaneous must remain tentative.
@@
Science is compatible with a form of pluralism that affirms the reality of higher-level causal powers (99). As scientists, we must use language carefully and remain grounded in physical mechanisms. The word emergence is highly used in the field of complex systems. While vague or mystical references to emergence may sound compelling, selling complexity science by mystifying emergence and invoking some spiritual dimensions is a great disservice (100). We seek to clarify emergence by framing it as a many-to-one coarse-graining process that retains predictive power while discarding most micro-level details. Within this framework, we differentiate between weak emergence, which is fully compatible with underlying dynamics, and the more philosophically problematic notion of strong emergence. Furthermore, we mention how measurable quantities—such as information flow, causal strength, order-parameter dynamics, and symmetry breaking—can transform emergence from mere rhetoric into testable science. Finally, as a concrete example in a more social setting, we mentioned recent work on herd immunity in structured contact networks where emergence can be rigorously defined, quantified, and linked to network geometry. Understanding the emergence of herd immunity highlights that epidemic control, like many collective phenomena, cannot be fully understood without considering the structure and correlation across social scales.
@@
AKR is grateful to Petter Holme and Yasser Roudi for their thoughtful discussions and acknowledges support from the Independent Research Fund Denmark (EliteForsk grant to Sune Lehmann), the Carlsberg Foundation (The Hope Project), and the Villum Foundation (NationScale Social networks). For this manuscript, grammar and style improvements were aided by AI language tools, under the author’s full supervision and approval.

### v3

@@ A New Ontology
The emergence of the giant component is foundational: the Internet stays functional despite random failures because its giant component remains intact (125). Likewise, in epidemic dynamics, if an infected person enters a social network and contacts someone in the giant component, the disease can potentially spread to nearly everyone after enough time. In this way, the onset of an epidemic reflects the same percolation principles that govern network connectivity (126, 127). We say that an outbreak occurs or an epidemic emerges when a significant number of individuals in a population become infected. The onset, size, and time scales of this emergence can be measured and modeled through the contagiousness of the disease and how the contact network is structured and changes over time (128). The emergence of large-scale connectivity is one example among many in networks. Furthermore, it is also easy to argue that communities emerge when networks evolve with local rules (129).
A mesoscale theory can clarify what aspects a good coarse-graining should focus on (130). For instance, when describing connectivity and spreading phenomena on networks at an intermediate level—specifically, how communities interact—it can capture most of the predictive signals while disregarding unnecessary details. At this scale, the dynamics are largely self-contained, with percolation and epidemic thresholds determined mainly by the pattern and strength of inter-module links (131). Identifying this mesoscale is an integral part of the theory itself (132, 133), as it is where compression, autonomy, and universality come together (130).
Emergence of Herd Immunity in a Population
@@
Vaccination is a key strategy for suppressing epidemics. Take measles, a highly contagious disease with a basic reproduction number of $R_{0}\sim 15$—meaning one case can infect around fifteen others early on in a fully susceptible population (134, 135). Before routine vaccination in the UK began in the late 1960s, outbreaks recurred every 2–3 years as the number of susceptible children quietly rose above the transmission threshold.
@@
Vaccination protects in two ways: directly, by immunizing individuals and removing them as nodes from the transmission network, and indirectly, as Fig. 3(a,b) shows, by forming a firewall (or firebreak (136)) and breaking the chains along which the infection might spread (137). In a vaccinated population, before the introduction of a new infection, individuals are either susceptible ($\mathcal{S}$) or immune ($\mathcal{R}$). The network representation of this population can always be rearranged to separate these two groups, with the interface formed by links connecting $\mathcal{S}$ and $\mathcal{R}$ nodes (Fig. 3f). The size of the immune set represents the direct benefit of vaccination, while the number of interface links serves as a proxy for indirect protection.
@@
Recent studies formalize this mechanism using bond percolation and message-passing on real and synthetic networks (140, 138). They track the dynamics of the susceptible giant component and the firefront density, $\rho_{\mathcal{SR}}$, across diverse mixing patterns. These analyses reveal a competing mechanism: targeting superspreaders or implementing acquaintance immunization increases collective immunity, while the spatial structure of the network can hinder it by localizing the protection, leaving remote, susceptible pockets vulnerable (140).
Herd immunity is a paradigmatic case of emergence in social networks: local immunizations combine nonlinearly to produce a global shield shaped by network structure. Recognizing this deepens our grasp of collective behavior and offers practical tools for designing more efficient interventions. Yet translating this insight into policy is non-trivial, given heterogeneous and mobile populations, vaccine profiles that differentially affect severity, infection, and carriage, and constraints on supply, waning, immune escape, uptake, equity, and ethics. Network-informed models should therefore complement—rather than replace—clinical priorities and socioeconomic evaluation, chiefly to enhance indirect protection where appropriate. The aim is disciplined decision support, not a prescriptive blueprint.
@@
In scientific practice, “emergence” almost always refers to weak emergence: a phenomenon grounded in the system’s components and interactions at the right scale, with no need for non-physical explanations. One can argue that a microscopic theory is incomplete or inapplicable to some macroscopic regimes—but to invoke strong emergence in domains like consciousness or social behavior (143, 144) amounts to introducing new causal principles not anchored in substrate dynamics, and thus steps outside the scope of scientific method (145). Therefore, if a macro-level description effectively explains a phenomenon—even more parsimoniously, with fewer micro-level details than a microscopic one—nothing beyond physicalism is at play. There have been, of course, historical movements that made strong-sounding claims about emergent phenomena without yielding testable or predictive frameworks (146, 147, 112, 148, 149, 150).
As causality connects events within the same scale, emergence connects descriptions across scales, revealing how collective patterns arise from local interactions. But the influence does not only run upward. Much like the walls of a container shape the motion of the particles inside, higher-level structures can constrain or guide the behavior of their components without violating the underlying laws (151). This interplay is often described as downward or top-down causation (152, 153, 154). Dominance hierarchies in animal groups illustrate this bidirectional linkage: dyadic encounters build rank, and that rank feeds back to channel individual aggression toward nearby rivals (151).
It is tempting to wonder whether complex aesthetic outcomes might also emerge from simple, knowable constraints—a question deeply resonant with the concept of emergence itself. If Poe’s self-portrait in The Philosophy of Composition (1846) is taken at face value, at least one kind of aesthetic complexity is not strongly emergent; it can be reconstructed from an explicit, step-wise recipe. Poe famously crafted the poem The Raven by first fixing a desired emotional effect—melancholy—and then deliberately selecting every element (theme, meter, and the refrain “Nevermore”) to achieve that goal (155). He insists good writing is methodical and analytical, explicitly rejecting artistic intuition. While many, including T. S. Eliot, argue that something ineffable lies beyond algorithmic methods (156), Poe’s essay reminds us that not all creativity is opaque to analysis—even if many artists and readers prefer it that way. Today, sentiment-analysis tools can map—and perhaps replicate—a narrative’s emotional arc (157, 158), sharpening our insight into how writers engineer intricate effects. However, until we can define such processes rigorously and run controlled experiments on well-defined creative variables, any final judgment about whether art is ultimately rule-bound or irreducibly spontaneous must remain tentative.
@@
In this work, we show the onset of emergence can be located through complementary lenses: in symmetry-breaking transitions, an order parameter turns on at a critical point while susceptibility and correlation length diverge; in RG terms, tuning the control parameter drives the flow to a scale-invariant fixed point so macroscopic behavior becomes largely insensitive to microscopic detail (84, 87, 88). We seek to clarify emergence by framing it as a many-to-one coarse-graining process that retains predictive power while discarding most micro-level details. Within this framework, we differentiate between the [weak] emergence, which is fully compatible with underlying dynamics, and the more philosophical notion of strong emergence. Finally, as a concrete example in a more social setting, we mentioned recent work on herd immunity in structured contact networks where emergence can be rigorously defined, quantified, and linked to network geometry. Understanding the emergence of herd immunity highlights that epidemic control, like many collective phenomena, cannot be fully understood without considering the structure and correlation across social scales.
Emergence is not a license to skip reason, nor a metaphysical extra cause (159). Higher-level patterns (from thermodynamic properties to life itself) can be real and explanatory while remaining grounded in physical laws (160). Any genuine emergent feature must still emerge from somewhere, meaning there is always an underlying mechanism, not beyond the standard physical interactions, producing it.(161). Science is compatible with a form of pluralism that affirms the reality of higher-level causal powers (162). As scientists, we must use language carefully and remain grounded in physical mechanisms. The word emergence is highly used in the field of Complex Systems. While vague or mystical references to emergence may sound compelling, selling complexity science by mystifying emergence and invoking some spiritual dimensions is a great disservice (163).
@@
AKR is grateful to Petter Holme, Yasser Roudi, Hossein Mahdaei, and Behrad Taghavi for all their conversations around this topic. This project was supported by the Independent Research Fund Denmark (EliteForsk grant to Sune Lehmann), the Carlsberg Foundation (The Hope Project), and the Villum Foundation (NationScale Social Networks). For this manuscript, grammar and style improvements were aided by AI language tools under the author’s full supervision and approval.